OpenStudy (anonymous):

a truck with 48 in. diameter wheels is travelling at 50mi/h a. find the angular speed of the wheels rad/min. b. how many revolutions per minute do the wheels make?

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

I'm not that good at these problems.. x*(

OpenStudy (anonymous):

i know the formula for this though lol

OpenStudy (anonymous):

for angular speed it should be \[\omega =\theta \div time\]

OpenStudy (anonymous):

Remember that radians are distance divided by the raidus.

OpenStudy (anonymous):

Do you know how to convert all the units?

OpenStudy (anonymous):

The radius will be \(24\) in.

OpenStudy (anonymous):

yes and the rotaions would be 50mi/ 60mins

OpenStudy (anonymous):

\[ \theta = \frac{x}{r}\\ \omega = \frac{v}{r} \]

OpenStudy (anonymous):

which one og the formulas would be used here for angular speed i thought it was \[\theta \div time\]

OpenStudy (anonymous):

You aren't given any time here, so that formula is of no use.

OpenStudy (anonymous):

You need to think about the relationship of everything.

OpenStudy (anonymous):

First of all, what is \(v\)?